OpenStudy (anonymous):

Factor (3x+2y)^3+(y-3)^3

OpenStudy (anonymous):

And i already know the formula, i just cant seem to get the correct answer

myininaya (myininaya):

Ok \[a^3+b^3=(a+b)(a^2-ab+b^2)\] So what is a and b here? a=(3x+2y) b=(y-3) Plug in and have fun

OpenStudy (anonymous):

like i said, i plug it in, and still get the wrong answer, i end up getting 9x^2+12xy+4y^2-3xy-9x-6y+2y^2+y^2-6y+9

myininaya (myininaya):

Well it says factor

myininaya (myininaya):

You didn't leave it in factored form

OpenStudy (anonymous):

(3x+2y+y-3)(9x^2+12xy+4y^2- 3xy-9x-6y+2y^2+y^2 -6y +9)

myininaya (myininaya):

(3x+2y)(y-3) 3x(y-3)+2y(y-3) 3xy-9x+2y^2-6y --- -(3x+2y)(y-3) -(3xy-9x+2y^2-6y) -3xy+9x-2y^2+6y

myininaya (myininaya):

You didn't distribute that -1 correctly

OpenStudy (anonymous):

what was the (3x+2y)(y-3), was that for the second part of the equation?

myininaya (myininaya):

The -ab part

OpenStudy (anonymous):

how did this (3x+2y)(y-3) go to this 3x(y-3)+2y(y-3)

myininaya (myininaya):

\[(a+b)(c+d)=a(c+d)+b(c+d)\]

myininaya (myininaya):

\[(a+b+e)(c+d)=a(c+d)+b(c+d)+e(c+d)\]

myininaya (myininaya):

\[(a+b+e+j)(c+d+e)=a(c+d+e)+b(c+d+e)+e(c+d+e)+j(c+d+e)\]

OpenStudy (anonymous):

so what factored out of the (3x+2y) to get (y-3)

myininaya (myininaya):

?

myininaya (myininaya):

I was just saying you didn't distribute the -1 in front of ab correctly

myininaya (myininaya):

do you understand what i'm saying?

myininaya (myininaya):

\[a^3+b^3=(a+b)(a^2-ab+b^2) \] a=3x+2y b=y-3

myininaya (myininaya):

See that our formula requires -ab?

myininaya (myininaya):

\[-ab=-(3x+2y)(y-3)\] \[=-[3x(y-3)+2y(y-3)]\] \[=-[3xy-9x+2y^2-6y]\] \[=-3xy+9x-2y^2+6y\]

myininaya (myininaya):

You have the a^2 and b^2 part right just the -ab part was wrong

myininaya (myininaya):

\[(3x+2y+y-3)(9x^2+12xy+4y^2-3xy+9x-2y^2+6y+y^2-6x+9)\] Simplify the insiders of each ( ) and you are done

OpenStudy (anonymous):

(3x+3y-3)(9x^2-9xy+3y^2-3x+6y+9)

myininaya (myininaya):

I see 9x-6x which is 3x not -3x

myininaya (myininaya):

and 12xy-3xy=9xy not -9xy

myininaya (myininaya):

Ok I think that looks right now you could actually factor more lol

OpenStudy (anonymous):

3(x+y-1)3(3x^2+3xy+y^2+x+2y+3)?

myininaya (myininaya):

Right and then 3*3=9

OpenStudy (anonymous):

so then 9(x+y-1)(3x^2+3xy+y^2+x+2y+3), so is that the final answer?

myininaya (myininaya):

Looks good to me

OpenStudy (anonymous):

I don't think there is suppose to be a 2y though

OpenStudy (anonymous):

I messed up somewhere, cuz the final answer in the book doesn't have 2y

myininaya (myininaya):

yeah the 6y-6y=0 so there is no term in the ( ) with just variable part y

myininaya (myininaya):

\[(3x+2y+y-3)(9x^2+12xy+4y^2-3xy+9x-2y^2+6y+y^2-6y+9) \] see in the second parathesis 6y-6y=0

myininaya (myininaya):

and also there is 9x

myininaya (myininaya):

\[(3x+3y-3)(9x^2+9xy+6y^2+9x+9)\]

myininaya (myininaya):

Now factor the 3's out from each ( )

OpenStudy (anonymous):

oh okay i see it